OpenStudy (anonymous):

how do you work a fraction like this 2-3/4

sam (.sam.):

\[2-\frac{3}{4}\] \[1\frac{4}{4}-\frac{3}{4}\] \[1\frac{4-3}{4}\] \[1\frac{1}{4}=\frac{5}{4}\]

OpenStudy (anonymous):

thanks